OpenStudy (happysunshine44):

Determine if the function has a maximum or minimum value, find that value and state the domain and range.

OpenStudy (happysunshine44):

\[f(x)=5x ^{2}\]

OpenStudy (happysunshine44):

@Mehek14

OpenStudy (learner):

When \(x=0\), you get \(f(x)=0\). When \(x=\pm1\), you get \(f(x)=(\pm 1)^2=1\). When \(x=\pm2\), you get \(f(x)=(\pm 2)^2=4\). When \(x=\pm3\), you get \(f(x)=(\pm 3)^2=9\). And so on .... \(\color{blue}{\rm Note~that~the~output~can~get~as~large~as~you~wish~-~forever, }\) \(\color{blue}{\rm (in~other~words,~there~isn't~a~maximum),~and~the~smallest }\) \(\color{blue}{{\rm value~you~may~get~for~}f(x){~\rm is~}f(x)=0,{\rm ~at~}x=0. }\)